# D-69120 Heidelberg, Germany

### v2

@@ 4 The radiation field
Typical sources of radiation are chosen to evaluate the RATs they cause by irradiating the grains. Among the different sources are two O-type stars labeled S1 and S2, respectively, with a flux of
@@
where $R$ is the stellar radius and $d$ is the distance from the star ($1\,\mathrm{pc}$). The emission of an active galactic nucleus (AGN) is modeled as
@@
following Giang et al. (2022) and Nenkova et al. (2008), where $C$ is a constant chosen such that the integration of $L^{\mathrm{AGN}}_{\lambda}$ equals a typical AGN bolometric luminosity of ${L_{\mathrm{bol}}=10^{10}L_{\odot}}$.
Finally, the spectral flux of the interstellar radiation field (ISRF) is parameterized as
@@
following Camps et al. (2015), based on Mathis et al. (1983) and Weingartner & Draine (2001). The applied parameters of the radiation sources are listed in Tab. 2 and the resulting spectral flux of the sources is shown in Fig. 2.
@@ 5 Optical and mechanical dust properties
To test the predictions of the RAT theory (LH07; Hoang et al. 2014) we consider silicates (SC) and carbonaceous (CA) grains, typical materials present in the ISM (see e.g. Draine 2003, for review). The optical properties of the materials are represented by the wavelength-dependent complex refractive index ${\bar{n}=n+ik}$. With $n$ being the real part describing refraction of light interacting with the dust grain and the imaginary part $k$ governs the light attenuation. For the SC take the (Astro)silicate model data as presented in Draine & Lee (1984); Laor & Draine (1993) and the optical properties of hydrogenated amorphous carbon are outlined in Jones (2012) 333 In detail: The refractive indices of the CA material is taken from http://vizier.cds.unistra.fr/viz-bin/VizieR?-source=J/A+A/542/A98 for $30\ \mathrm{nm}$ hydrocarbon particles with $E_{g}=2.67\,\mathrm{eV}$ and the SC-data is listed as eps_Sil in https://www.astro.princeton.edu/~draine/dust/dust.diel.html. Ice mantles were not considered here, as high intensity radiation environments would not allow them. Assuming condensation of all ice species at a temperature above $150\,\mathrm{K}$ (Zhang et al. 2015) and comparing with dust temperatures in Tab. 3, we would expect ice-free grains for the AGN source and for CA grains around the stellar sources. The ice-free assumption may therefore generally not be valid for the ISRF and SC grains around the stellar sources. Additionally comparing with these dust temperatures assumes there is no significant difference in dust temperature when a icy surface is considered instead of the silicate/carbonaceous surface.
The radiative torque efficiency ${Q}_{\Gamma}$ is calculated using the DDA code DDSCAT (version 7.3.3, see Draine & Flatau 1994, 2008; Flatau & Draine 2012). The grain aggregates are initialized and tabulated as outlined in Sect. 2. To create input shape-files for DDSCAT simulations, the volume enclosed by the monomers is filled with a cubic 3D-grid, each grid point representing a distinct dipole, with properties determined by the complex refractive index. DDSCAT calculations of ${Q}_{\Gamma}$ are performed in the external coordinate system ${\{\hat{e}_{1},\hat{e}_{2},\hat{e}_{3}\}}$, of incoming radiation
@@
where $\hat{e}_{\mathrm{1}}$ is anti-parallel to the wave-vector of incoming light. For simplicity, we assume that the grain is aligned with the direction of radiation i.e. $\Theta=0^{\circ}$. Subsequently, the direction of radiation and rotation coincide ($\hat{e}_{\mathrm{1}}\ ||\ \hat{a}_{\mathrm{1}}$) and the RAT efficiency factor becomes ${Q}_{\mathrm{\Gamma}}\equiv Q_{\mathrm{e1}}\left(\Theta\right)\hat{e}_{\mathrm{1}}$. Since RATs are most efficient for $\Theta=0^{\circ}$ the data presented in this paper represents merely an upper limits for the modeling of the spin-up process of ballistic aggregates. However, focusing on the special case of $\Theta=0^{\circ}$ does not impact our comparison with previous attempts of modeling the efficiency of RATs. Simulation results are averaged over 21 equidistant values of the rotation angle of the grain around $\hat{a}_{1}$. The results are computed for 50 wavelengths between $0.05\,\mu\mathrm{m}$ and $150\,\mu\mathrm{m}$ within equidistant bins in log-space. Each combination of material, monomer-density, size, and shape (see Sect. 2) is simulated with five different resolutions i.e. the separation of two nearest dipoles. Those are combined using an average weighted by dipole resolution. In addition to the RAT efficiency factor ${Q}_{\mathrm{\Gamma}}$, DDSCAT also provides the efficiency of absorption ${Q}_{\mathrm{abs}}$ required to evaluate the photon-drag-timescale $\tau_{\mathrm{IR}}$ and the dust temperature $T_{\mathrm{dust}}$. The gas-torque-efficiency $Q_{\mathrm{gas}}$ needed to evaluate the gas-drag-timescale $\tau_{\mathrm{gas}}$ is calculated with a Monte Carlo (MC) based gas-dust collision approach as outlined in Reissl et al. (2023). The corresponding gas parameters [^4] We note that the gas number density $n_{\mathrm{gas}}$ is lower than in a typical AGN torus. However, we can position the dust grain outside the torus, selecting a position in which $n_{\mathrm{gas}}$ is similar to the ISM, e.g. at higher inclinations from the plane of the torus. are listed in Tab. 1.
@@
Figure 3: Left panel: The drag timescales over the effective radius $a_{\mathrm{eff}}$ dependent on grain model (BA, BAM1, BAM2) and different gas densities $n_{\mathrm{gas}}$ for all grain materials. Here, the gas temperature is at a constant value of $T_{\mathrm{gas}}=100\ \mathrm{K}$. Center panel: The same as the left panel but for the IR drag timescale for the CA grain material, for different dust temperatures $T_{\mathrm{dust}}$. Right panel: The same as center panel but for the SC grain material.
@@ 6.1 Drag timescales
Resulting timescales for the gas-drag and the photon-drag, respectively, taking different gas densities and dust temperatures are shown in Fig. 3. The gas-drag-timescale $\tau_{\mathrm{gas}}$ is presented in the left panel as an average of SC and CA materials. We emphasize that the gas-drag-efficiency $Q_{\mathrm{gas}}$ is only dependent on grain shape in our modeling and is independent of material. We see that denser gas equals a smaller $\tau_{\mathrm{gas}}$ in agreement with the literature (see e.g. Draine & Weingartner 1997 and Reissl et al. 2023). The dependence on monomer-density is marginal, with the BA grains having a stronger gas-drag than the BAM1 and BAM2 grains. This is due to the dependence of $Q_{\mathrm{gas}}$ on the surface area.
@@
In detail, the BA grains present a larger surface area for interaction with gas particles. Following Eq.5, the expected size dependence of $\tau_{\mathrm{gas}}$ would be linear ($I_{\mathrm{max}}\propto a_{\mathrm{eff}}^{5}$). However it is almost constant above $100\,\mathrm{nm}$, which is due to $Q_{\mathrm{gas}}$ increasing linearly with size. In principle the gas drag efficiency $Q_{\mathrm{gas}}$ should only depend on the shape of the grain and not the size. That $Q_{\mathrm{gas}}$ nonetheless increases with size may be explained by $a_{\mathrm{eff}}$ not being well suited to describe grain size in this situation. The calculation of $Q_{\mathrm{gas}}$ uses the simulated gas-drag-torque, which depends on the angular momentum lost by the grain when a gas particle leaves the surface. This in turn scales linearly with the distance from the axis of rotation of the grain from where the gas particle is ejected. For porous grains this distance is often larger than $a_{\mathrm{eff}}$. The torque simulated in this way contains the information about the true shape of the grain, however, in the formula for $Q_{\mathrm{gas}}$ the size $a_{\mathrm{eff}}$ is used, which creates the size dependence of $Q_{\mathrm{gas}}$.
@@
The photon-drag-timescale is shown in the middle for CA grains and on the right for SC grains in Fig. 3. As the efficiency of absorption $Q_{\mathrm{abs}}$ is overall larger for SC, more photons are absorbed and emitted causing a stronger photon-drag. The number of emitted photons also depends on dust temperature, with hotter dust experiencing stronger drag (see also Tab. 3). The dependence on monomer-density is marginal, and is related to the fact that the moment of inertia is larger in the case of BA grains than for the more compact BAM2 grains given the same effective radius (see Appendix A in Reissl et al. 2023).
@@
Figure 4: Torque efficiency $Q_{\Gamma}$ as a function of wavelength $\lambda$ for an exemplary CA/BA grain of size $a_{\mathrm{eff}}=200\,\mathrm{nm}$. The spectra of the two stellar sources S1 and S2, respectively, are shown in blue, normalized by their maximum value. We emphasize that this particular grain has two changes in sign resulting in a reduced net-torque over wavelengths.
@@
Figure 5: Average torque efficiency $|Q_{\Gamma}|$ over wavelengths for CA (left) and SC (right) BA-aggregates of different grain sizes $a_{\mathrm{eff}}$ (color coded) in comparison with the real part $n$ (dotted line) and imaginary part $k$ (dashed line) of the corresponding refractive indices. We note that characteristic features of $Q_{\Gamma}$ follow the shape of $k$.
@@
Figure 6: The same as Fig. 5 but for the average torque efficiency $Q_{\Gamma}$ over $\lambda/a_{\mathrm{eff}}$ in comparison with the conical power-law parametrization (red dotted line) as suggested in LH07.
@@
Figure 7: Two-power-law parametrization shown as a red line for CA (left) and SC (right) BA-grains in comparison with the full set of the numerical DDSCAT simulations (black dots) for grains of size $a_{\mathrm{eff}}=200\ \mathrm{nm}$ where the absolute value of the torque efficiency $|Q_{\Gamma}|$ are plotted as a function of $\lambda/a_{\mathrm{eff}}$. The constant part $c$ is fitted for each grain size up to a value of $\lambda/a_{\mathrm{eff}}=3$, the steeper power-law is fitted as a linear function to the logarithm of the data for all sizes combined in the interval $103<\lambda/a_{\mathrm{eff}}<$ 10, same for the shallow power-law in the interval $\lambda/a_{\mathrm{eff}}>20$. Finally, the two intersection points $a$ and $b$ are calculated and used to create a continuous parametrization for $|Q_{\Gamma}|$ (see Eq. 13).
@@ 6.2 Torque efficiency
As we assume that the grains principal axis $\hat{a}_{1}$ is parallel to $\hat{e}_{1}$, the direction of incoming radiation, only the $Q_{1}$ component needs to be evaluated to get the total torque efficiency $Q_{\Gamma}$.
@@
Results of the efficiency $Q_{\Gamma}$ calculated with DDSCAT may possess positive (negative) values dependent on wavelength (see Fig. 4) corresponding to a direction of the angular velocity ${\omega}_{\mathrm{RAT}}$ being parallel (anti-parallel) to $\hat{e}_{1}$. Hence, different wavelength regimes create opposing torques on the grain aggregates. Subsequently, taking the sign change into account may considerably diminishing the net RAT for individual aggregates. In order to evaluate the average efficiency for different grain aggregates we both evaluate $\Gamma_{\mathrm{RAT}}$ over $\lambda$ (see Eq. 1) for each grain individually including canceling effects of positive and negative spectral contribution of $Q_{\Gamma}(\lambda)$ as well as calculate $\Gamma_{\mathrm{RAT}}$ using the absolute value $|Q_{\Gamma}(\lambda)|$, which ignores the opposing spectral contributions to the net RAT. We find that averaging over the DDSCAT data with the absolute value of $|Q_{\Gamma}|$ results only in marginally higher RATs compared to taking the sign of $Q_{\Gamma}$ into account. We emphasize that the difference is insignificant compared to the variation between individual grain shapes. Hence, we utilize only the absolute value $|Q_{\Gamma}|$ in the following analysis, to make it possible to combine data of different grain shapes.
@@
The resulting dependence of the magnitude of $Q_{\Gamma}$ on wavelength for CA/BA and SC/BA grains is presented in Fig. 5. The values of $Q_{\Gamma}$ are compared with the refractive index of the corresponding grain material. We report a feature for SC at $\lambda>8\,\mu\mathrm{m}$, that follow the form of the imaginary part $k$ of the refractive index. This feature of SC is also noted in previous studies (see e.g. LH07).
6.3 Parametrization for the RAT efficiency
@@
Since RATs depend on the wavelength $\lambda$ relative to the grain radius $a_{\mathrm{eff}}$ it is convenient to evaluate $|Q_{\Gamma}|$ as a function of $\lambda/a_{\mathrm{eff}}$ (see Fig. 6) instead of $\lambda$ alone. Here, the slopes of the different grain sizes overlap tightly, showing the self-similar nature that allowed the canonical parametrization (Eq. 2) for solid grains in the first place. Our results using grain aggregates shows a comparable constant part for $\lambda/a_{\mathrm{eff}}<1.8$, but do not follow the overall power-law parametrization for $\lambda/a_{\mathrm{eff}}>1.8$. Instead both materials SC and CA, reveal a steeper power-law that overlaps for all sizes from $\lambda/a_{\mathrm{eff}}\approx 3$ and then goes over into a shallower power-law for $\lambda/a_{\mathrm{eff}}\gtrsim 10$. A similar behavior of a broken power-law can be seen in previous studies e.g. in LH07 or for gaussian dust grains in Herranen et al. (2019), but is much more apparent here. Consequently, we suggest a modified parametrization for ballistic grain aggregates with two connected consecutive power-laws. Least square fits for the constant part (called $c$ in this study) and the two power-laws are performed for all data points of the grains sharing the same material and monomer-density. The fits of $c$ are performed for individual grain sizes in the interval $\lambda/a_{\mathrm{eff}}<3$. All sizes are fitted jointly for the first and second power-law with exponents $q_{1}$ and $q_{2}$ in the intervals $123<(\lambda/a_{\mathrm{eff}})<$ 12 and $(\lambda/a_{\mathrm{eff}})>20$. In practice the power-laws are fitted as linear functions to the logarithm of the data.
@@
In order to combine the three regions conveniently into a continuous curve the following definition was chosen:
@@
Here, the individual fits are joined by calculating the intersection points $a$ and $b$, where $a$ is the intersection of $c$ with the steep power-law and $b$ is the one of both power-laws. Therefore like $c$, there will be a different value of $a$ for each $a_{\mathrm{eff}}$. The resulting parametrization is shown in Fig. 7 for the exemplary data set of carbon BA-grains of size $200\,\mathrm{nm}$. All resulting fit parameters of Eq. 13 are given in Appendix A.
Larger grains interact more efficiently with the radiation field resulting in a higher value $c$ as can be seen in Fig. 6. The BAM1 and BAM2 grain models have comparable curves for $|Q_{\Gamma}|$, where the values differ most in the constant part on a scale similar to the difference between two grain sizes. The more porous the grain the larger $c$. Differences between the two materials are the values of $c$, which are larger for SC. Both power-laws are steeper for CA, most likely due to interference by the feature in the SC material.

### v3

@@ 4 The radiation field
Typical sources of radiation were chosen to evaluate the RATs they cause by irradiating the grains. The different sources included two O-type stars labeled S1 and S2, with a flux of
@@
where $R$ is the stellar radius, and $d$ is the distance from the star ($1\,\mathrm{pc}$). The emission of an active galactic nucleus (AGN) was modeled as
@@
following Giang et al. (2022) and Nenkova et al. (2008), where $C$ is a constant chosen such that the integration of $L^{\mathrm{AGN}}_{\lambda}$ equals the typical AGN bolometric luminosity of ${L_{\mathrm{bol}}=10^{10}L_{\odot}}$.
Finally, the spectral flux of the interstellar radiation field (ISRF) was parameterized as
@@
following Camps et al. (2015), which is based on Mathis et al. (1983) and Weingartner & Draine (2001). The applied parameters of the radiation sources are listed in Tab. 2, and the resulting spectral flux of the sources is shown in Fig. 2.
@@ 5 Optical and mechanical dust properties
To test the predictions of the RAT theory (LH07; Hoang et al. 2014), we considered silicates (SC) and carbonaceous (CA) grains. These materials are typically present in the ISM (see e.g. Draine 2003, for review). The optical properties of the materials are represented by the wavelength-dependent complex refractive index ${\bar{n}=n+ik}$, with $n$ being the real part describing the refraction of light that interacts with the dust grain, and the imaginary part $k$ governs the light attenuation. For the SC, we took the (Astro)silicate model data as presented in Draine & Lee (1984); Laor & Draine (1993), and the optical properties of hydrogenated amorphous carbon were outlined in Jones (2012) 333 In detail: The refractive indices of the CA material were taken from http://vizier.cds.unistra.fr/viz-bin/VizieR?-source=J/A+A/542/A98 for $30\ \mathrm{nm}$ hydrocarbon particles with $E_{g}=2.67\,\mathrm{eV}$, and the SC data are listed as eps_Sil in https://www.astro.princeton.edu/~draine/dust/dust.diel.html. Ice mantles were not considered here because high-intensity radiation environments would not allow them. Assuming condensation of all ice species at a temperature above $150\,\mathrm{K}$ (Zhang et al. 2015) and comparing with dust temperatures in Tab. 3, we would expect CA grains to be ice-free for the stellar sources as well as both CA and SC grains for the AGN source. The ice-free assumption may therefore generally not be valid for the ISRF and SC grains around the stellar sources. Additionally, a comparison with these dust temperatures assumes that there is no significant difference in the dust temperature when an icy surface is considered instead of the silicate or carbonaceous surface.
The radiative torque efficiency ${Q}_{\Gamma}$ was calculated using the DDA code DDSCAT (version 7.3.3; see Draine & Flatau 1994, 2008; Flatau & Draine 2012). The grain aggregates were initialized and tabulated as outlined in Sect. 2. To create input shape files for the DDSCAT simulations, we filled the volume enclosed by the monomers with a cubic 3D grid, each grid point representing a distinct dipole, with properties determined by the complex refractive index. DDSCAT calculations of ${Q}_{\Gamma}$ were performed in the external coordinate system ${\{\hat{e}_{1},\hat{e}_{2},\hat{e}_{3}\}}$ of incoming radiation,
@@
where $\hat{e}_{\mathrm{1}}$ is antiparallel to the wave vector of incoming light. For simplicity, we assumed that the grain was aligned with the direction of radiation, that is, $\Theta=0^{\circ}$. Subsequently, the direction of radiation and rotation coincide ($\hat{e}_{\mathrm{1}}\ ||\ \hat{a}_{\mathrm{1}}$), and the RAT efficiency factor becomes ${Q}_{\mathrm{\Gamma}}\equiv Q_{\mathrm{e1}}\left(\Theta\right)\hat{e}_{\mathrm{1}}$. Since RATs are most efficient for $\Theta=0^{\circ}$, the data presented in this paper merely represent an upper limit for the modeling of the spin-up process of ballistic aggregates. However, the focus on the special case of $\Theta=0^{\circ}$ does not impact our comparison with previous attempts of modeling the efficiency of RATs. The simulation results were averaged over 21 equidistant values of the rotation angle of the grain around $\hat{a}_{1}$. The results were computed for 50 wavelengths between $0.05\,\mu\mathrm{m}$ and $150\,\mu\mathrm{m}$ within equidistant bins in log-space. Each combination of material, monomer density, size, and shape (see Sect. 2) was simulated with five different resolutions, that is, the separation of two nearest dipoles. They were combined using an average weighted by the dipole resolution. In addition to the RAT efficiency factor ${Q}_{\mathrm{\Gamma}}$, DDSCAT also provides the efficiency of the absorption ${Q}_{\mathrm{abs}}$ that is required to evaluate the photon-drag timescale $\tau_{\mathrm{IR}}$ and the dust temperature $T_{\mathrm{dust}}$. The gas-torque efficiency $Q_{\mathrm{gas}}$ needed to evaluate the gas-drag timescale $\tau_{\mathrm{gas}}$ was calculated with a Monte Carlo-based (MC) gas-dust collision approach as outlined in Reissl et al. (2023). The corresponding gas parameters [^4] We note that the gas number density $n_{\mathrm{gas}}$ is lower than in a typical AGN torus. However, we can position the dust grain outside the torus, selecting a position in which $n_{\mathrm{gas}}$ is similar to the ISM, e.g., at higher inclinations from the plane of the torus. are listed in Tab. 1.
@@
Figure 3: Drag timescales over the effective radius $a_{\mathrm{eff}}$ dependent on the grain model (BA, BAM1, or BAM2). Left panel: Gas drag timescale for different gas densities $n_{\mathrm{gas}}$ for all grain materials. The gas temperature is at a constant value of $T_{\mathrm{gas}}=100\ \mathrm{K}$. Center panel: IR drag timescale for the CA grain material for different dust temperatures $T_{\mathrm{dust}}$. Right panel: Same as the center panel, but for the SC grain material.
@@ 6.1 Drag timescales
We show the resulting timescales for the gas and photon drag for different gas densities and dust temperatures in Fig. 3. The gas-drag timescale $\tau_{\mathrm{gas}}$ is presented in the left panel as an average of SC and CA materials. We emphasize that the gas-drag efficiency $Q_{\mathrm{gas}}$ is only dependent on the grain shape in our modeling and is independent of the material. Denser gas equals a lower $\tau_{\mathrm{gas}}$, in agreement with the literature (see, e.g., Draine & Weingartner 1997 and Reissl et al. 2023). The dependence on the monomer density is marginal, with the BA grains having a stronger gas drag than the BAM1 and BAM2 grains. This is due to the dependence of $Q_{\mathrm{gas}}$ on the surface area.
@@
In detail, the BA grains present a larger surface area for interaction with gas particles. Following Eq.5, the expected size dependence of $\tau_{\mathrm{gas}}$ would be linear ($I_{\mathrm{max}}\propto a_{\mathrm{eff}}^{5}$). It is almost constant above $100\,\mathrm{nm}$, however, which is due to the linear increase in $Q_{\mathrm{gas}}$ with size. In principle, the gas drag efficiency $Q_{\mathrm{gas}}$ should only depend on the shape of the grain and not on the size. $Q_{\mathrm{gas}}$ nonetheless increases with size, and the reason for this may be that $a_{\mathrm{eff}}$ is poorly suited to describe the grain size in this situation. The calculation of $Q_{\mathrm{gas}}$ used the simulated gas-drag-torque, which depends on the angular momentum lost by the grain when a gas particle leaves the surface. This in turn scales linearly with the distance from the axis of rotation of the grain from which the gas particle is ejected. For porous grains, this distance is often larger than $a_{\mathrm{eff}}$. The torque simulated in this way contains information about the true shape of the grain, but in the formula for $Q_{\mathrm{gas}}$ , the size $a_{\mathrm{eff}}$ is used, which creates the size dependence of $Q_{\mathrm{gas}}$.
@@
The photon-drag timescale is shown in the middle for CA grains and on the right for SC grains in Fig. 3. As the efficiency of absorption $Q_{\mathrm{abs}}$ is higher overall for SC, more photons are absorbed and emitted, which causes a stronger photon drag. The number of emitted photons also depends on the dust temperature, with hotter dust experiencing stronger drag (see also Tab. 3). The dependence on the monomer density is marginal, and this is related to the fact that the moment of inertia is higher in the case of BA grains than for the more compact BAM2 grains for the same effective radius (see Appendix A in Reissl et al. 2023).
@@
Figure 4: Torque efficiency $Q_{\Gamma}$ as a function of wavelength $\lambda$ for an exemplary CA BA grain of size $a_{\mathrm{eff}}=200\,\mathrm{nm}$. The spectra of the two stellar sources S1 and S2 are shown in blue, normalized by their maximum value. We emphasize that this particular grain has two changes in sign, resulting in a reduced net torque over wavelengths.
@@
Figure 5: Average torque efficiency $|Q_{\Gamma}|$ over wavelengths for CA (left) and SC (right) BA aggregates of different grain sizes $a_{\mathrm{eff}}$ (color-coded) in comparison with the real part $n$ (dotted line) and imaginary part $k$ (dashed line) of the corresponding refractive indices. We note that the characteristic features of $Q_{\Gamma}$ follow the shape of $k$.
@@
Figure 6: Same as Fig. 5, but for the average torque efficiency $Q_{\Gamma}$ over $\lambda/a_{\mathrm{eff}}$ in comparison with the conical power-law parameterization (dotted red line), as suggested in LH07.
@@
Figure 7: Two-power-law parameterization shown as a red line for CA (left) and SC (right) BA grains in comparison with the full set of the numerical DDSCAT simulations (black dots) for grains of size $a_{\mathrm{eff}}=200\ \mathrm{nm}$, where the absolute value of the torque efficiency $|Q_{\Gamma}|$ is plotted as a function of $\lambda/a_{\mathrm{eff}}$. The constant part $c$ is fitted for each grain size up to a value of $\lambda/a_{\mathrm{eff}}=3$, the steeper power-law is fit as a linear function to the logarithm of the data for all sizes combined in the interval $103<\lambda/a_{\mathrm{eff}}<$ 10, and the same for the shallow power law in the interval $\lambda/a_{\mathrm{eff}}>20$. Finally, the two intersection points $a$ and $b$ were calculated and used to create a continuous parameterization for $|Q_{\Gamma}|$ (see Eq. 13).
@@ 6.2 Torque efficiency
As we assumed that the principal grain axis $\hat{a}_{1}$ is parallel to $\hat{e}_{1}$, the direction of the incoming radiation, only the $Q_{1}$ component needs to be evaluated to obtain the total torque efficiency $Q_{\Gamma}$.
@@
The results of the efficiency $Q_{\Gamma}$ calculated with DDSCAT may possess positive (negative) values depending on the wavelength (see Fig. 4), corresponding to a direction of the angular velocity ${\omega}_{\mathrm{RAT}}$ that is parallel (antiparallel) to $\hat{e}_{1}$. Hence, different wavelength regimes create opposing torques on the grain aggregates. When the sign change is subsequently taken into account, the net RAT for individual aggregates may considerably lower. In order to evaluate the average efficiency for different grain aggregates, we evaluated $\Gamma_{\mathrm{RAT}}$ over $\lambda$ (see Eq. 1) for each grain individually, including canceling effects of positive and negative spectral contribution of $Q_{\Gamma}(\lambda)$, and we also calculated $\Gamma_{\mathrm{RAT}}$ using the absolute value $|Q_{\Gamma}(\lambda)|$, which ignores the opposing spectral contributions to the net RAT. We find that averaging over the DDSCAT data with the absolute value of $|Q_{\Gamma}|$ resulted in only marginally higher RATs compared to when the sign of $Q_{\Gamma}$ was taken into account. We emphasize that the difference is insignificant compared to the variation between individual grain shapes. Hence, we only used the absolute value $|Q_{\Gamma}|$ in the following analysis so that data of different grain shapes can be combined.
@@
The resulting dependence of the magnitude of $Q_{\Gamma}$ on the wavelength for CA BA and SC BA grains is presented in Fig. 5. The values of $Q_{\Gamma}$ are compared with the refractive index of the corresponding grain material. We report a feature for SC at $\lambda>8\,\mu\mathrm{m}$ that follows the form of the imaginary part $k$ of the refractive index. This feature of SC was also noted in previous studies (see, e.g., LH07).
6.3 Parameterization for the RAT efficiency
@@
Since RATs depend on the wavelength $\lambda$ relative to the grain radius $a_{\mathrm{eff}}$, it is convenient to evaluate $|Q_{\Gamma}|$ as a function of $\lambda/a_{\mathrm{eff}}$ (see Fig. 6) instead of $\lambda$ alone. Here, the slopes of the different grain sizes overlap tightly, showing the self-similar nature that allowed the canonical parameterization (Eq. 2) for solid grains in the first place. Our results using grain aggregates show a comparable constant part for $\lambda/a_{\mathrm{eff}}<1.8$, but do not follow the overall power-law parameterization for $\lambda/a_{\mathrm{eff}}>1.8$. Instead, both materials SC and CA reveal a steeper power law that overlaps for all sizes from $\lambda/a_{\mathrm{eff}}\approx 3$ and then evolves into a shallower power law for $\lambda/a_{\mathrm{eff}}\gtrsim 10$. A similar behavior of a broken power law was reported in previous studies, for instance, in LH07, or for Gaussian dust grains in Herranen et al. (2019), but is much more apparent here. Consequently, we suggest a modified parameterization for ballistic grain aggregates with two connected consecutive power laws. We performed least-square fits for the constant part (called $c$ in this study) and the two power laws for all data points of the grains that shared the same material and monomer density. The fits of $c$ were performed for individual grain sizes in the interval $\lambda/a_{\mathrm{eff}}<3$. All sizes were fit jointly for the first and second power law with exponents $q_{1}$ and $q_{2}$ in the intervals $123<(\lambda/a_{\mathrm{eff}})<$ 12 and $(\lambda/a_{\mathrm{eff}})>20$. In practice, the power laws were fit as linear functions to the logarithm of the data.
@@
In order to conveniently combine the three regions into a continuous curve, we chose the following definition:
@@
We joined the individual fits by calculating the intersection points $a$ and $b$, where $a$ is the intersection of $c$ with the steep power law, and $b$ is the intersection of both power laws. Therefore, like $c$, there will be a different value of $a$ for each $a_{\mathrm{eff}}$. The resulting parameterization is shown in Fig. 7 for the exemplary data set of carbon BA grains of size $200\,\mathrm{nm}$. All resulting fit parameters of Eq. 13 are given in Appendix B.
Larger grains interact more efficiently with the radiation field, which results in a higher value $c$, as shown in Fig. 6. The BAM1 and BAM2 grain models have comparable curves for $|Q_{\Gamma}|$, where the values differ most in the constant part on a scale that is similar to the difference between two grain sizes. The more porous the grain, the larger $c$. The two materials differ in their values of $c$, which are higher for SC. Both power laws are steeper for CA, most likely due to interference by the feature in the SC material.